OpenStudy (anonymous):

Find the height of the pole if the shadow of a boy 5 ft. 6 in. tall is 2 ft. and the shadow of the pole is 2 yards. 192 inches 195 inches 198 inches

OpenStudy (anonymous):

5 ft. 6 in. = 66 in. 2 ft. = 24 in. 2 yds. = 72 in. 66/24 = x/72 = 66/1 = x/3 x = 66(3) = 198 in.

OpenStudy (campbell_st):

change all the measurements to inches... draw a diagram then use ratios of corresponding sides in similar triangles  then \[\frac{h}{66} = \frac{72}{24}\] just solve for h